# Universal Omni-Language Models

### v2

@@ 4 Experiment Settings
Gemini-1.5-Pro 47.56% 38.53% 34.68%
Reka-core-20240501 36.10% 35.07% 34.39%
Table 3: OLM Baselines Overall Results Grouped by Audio Type and Task Category. The accuracy calculated by different audio types are at the first three columns and the accuracy accross tasks categories are placed at the rest columns.
@@
Gemini-1.5-Pro 47.90% 46.36% 48.11%
Reka-core-20240501 38.06% 29.81% 37.74%
Model Action & Activity Story Description Plot Inference Object Identification & Description Contextual & Environmental Identity & Relationship Text & Symbols Count & Quantity
@@
Table 4: Results on Textual Audio Approximation Experiments. All the audios are represented in text transcript. The results are divided into groups of vision-language models and omni-models. We use the text transcript to approximate the audios in this setting.
@@
GPT4-o (0513) 57.62% 46.58% 43.73%
GPT4-o (0806) 53.46% 48.25% 38.35%
GPT4-o-mini 48.90% 39.58% 35.33%
Gemini-1.5-Pro 49.91% 31.79% 34.68%
Reka-core-20240501 48.73% 36.33% 34.39%
Claude-3.5-Sonnet 59.46% 37.48% 45.18%
GPT-4V-Preview 48.19% 44.75% 37.33%
GPT-4V-0409 42.49% 46.06% 37.72%
@@
LTU (7B) 23.38% 26.88% 23.29% 25.68% 19.25% 16.98%
LTU-AS (7B) 28.90% 25.31% 27.06% 29.31% 29.81% 23.58%
Mu-LLaMA (7B) 23.73% 25.74% 24.52% 22.70% 26.04% 25.47%
MusiLingo-long-v1 29.77% 25.22% 27.41% 28.79% 30.57% 34.91%
audio-SALMONN (13B) 26.18% 23.12% 30.39% 26.85% 24.91% 24.53%
Qwen-Audio-Chat 29.95% 28.72% 28.28% 27.89% 32.83% 37.74%
Qwen2-Audio-7B-Instruct 43.35% 37.72% 39.67% 43.19% 44.91% 40.57%
Audio-Flamingo (1.3B) 26.80% 25.74% 24.87% 25.29% 30.94% 27.36%
Gemini-1.5-Pro 43.93% 38.53% 30.21% 46.17% 39.62% 52.83%
Reka-core-20240501 34.57% 35.07% 32.94% 35.09% 33.21% 34.29%
@@
5 Findings
5.1 Results on Omni-Language Models
Despite overall poor performance, open-source baselines generally exhibit higher accuracy on speech audio, indicating a potential bias towards speech data. In contrast, Gemini-1.5-Pro and Reka-core-20240501, the two available proprietary models evaluated in this tri-modal setting, shows more promising results. Regarding the scores across audio types, the Gemini-1.5-Pro shows a more balanced performances while Reka-core-20240501 showing a lag on modeling the sound events. Moreover, the comparison of Gemini-1.5-Pro’s performance across full input context and ablated settings (image-text and audio-text) suggests that it effectively leverages information from all modalities to enhance its reasoning capabilities. While it demonstrates superior overall performance and balanced accuracy across audio types compared to open-source alternatives, its accuracy remains below 50%.
Breakdown Results.
We present the breakdown of the performance of open-source omni-language model baselines across different audio types and task categories in the OmniBench evaluation. The results reveal inconsistent performance patterns across audio types, with some models showing higher accuracy on sound events or music compared to speech. Across task categories, models tend to perform better on object identification and description tasks, while struggling with more reasoning tasks such as plot inference and story description. These findings highlight the current limitations of OLMs in integrating information across modalities.
5.2 Textual Approximation on Images and Audios
As the absence of effective OLMs on the OmniBench, we further introduce the text alternatives of the images ($I^{\prime}$) and audios ($A^{\prime}$) to embrace more dual-modal image-text or audio-text language models to analyze the research progress of the topic. Using the textual approximations of the multimodal contexts, we evaluate the baselines and put the results using audio transcripts, image captions, and both of them in Table 4, Table 5 and Table 6 correspondingly.
@@ 5.2 Textual Approximation on Images and Audios
We select the UnifiedIO-2 series OLMs to conduct the textual approximation experiments due to their relatively robust performances in the vanilla OmniBench evaluation setting suggested in Table 2. Compared with the vanilla setting, all three UnifiedIO-2 models show performance gains, averagely at $6.42\%$, in the audio replacement setting and average performance drops in the replaced-image ($1.87\%$) and both-repaced settings ($0.12\%$). This indicates the shortcoming of existing OLMs on modeling the audio on the one hand, and there is potentially more noise in the generated image captions compared to the human-written audio transcripts on the other hand.
@@
In the setting of using text as the alternatives of audios and images, the VLMs show ALMs a generally better results (Table 4 vs Table 5). This could be caused by : 1) more available research resources including datasets and methods are allocated to image-text models; 2) the general audio are naturally harder to annotate. If the textual alternatives of images and audios could regarded as in equal quality, it seems to be easier for the future researcher to train an omni-language model from exisiting VLMs rather than ALMs.
@@
LTU (7B) 22.77% 23.91% 22.59%
LTU-AS (7B) 29.16% 26.01% 26.09%
Mu-LLaMA (7B) 26.27% 25.74% 24.52%
MusiLIngo-long-v1 (7B) 28.46% 30.91% 27.41%
audio-SALMONN-13B 28.46% 27.93% 30.39%
Qwen-Audio-Chat 31.87% 34.15% 28.28%
Qwen2-Audio-7B-Instruct 47.02% 39.05% 39.67%
Audio-Flamingo (1.3B) 23.99% 25.92% 24.87%
@@
Cambrian-13B 43.08% 231.96% 29.16%
GPT4-o (0513) 60.60% 46.58% 38.79%
GPT4-o (0806) 56.76% 48.25% 36.69%
GPT4-o-mini 51.41% 39.58% 33.86%
Gemini-1.5-Pro 45.27% 31.79% 30.21%
Claude-3.5-Sonnet 57.44% 37.48% 41.33%
@@
Pure Textual Evaluation.
The results in Table 6 reveal interesting patterns in model performance when both image and audio inputs are replaced with textual descriptions. Notably, the majority of models demonstrate improved accuracy when processing textual representations of multimodal data compared to their performance on either image captions or audio transcripts alone. This suggests that these models are better equipped to integrate information from multiple textual sources rather than handling raw multimodal inputs. For instance, Qwen2-Audio-7B-Instruct shows a significant jump in accuracy from 39.05% (audio transcript only) and 39.67% (image caption only) to 47.02% when given both textual representations. Similarly, proprietary models like GPT4-o and Claude-3.5-Sonnet exhibit substantial gains, with GPT4-o (0513) achieving an impressive 60.60% accuracy in the pure textual setting.
6 Conclusion and Future Study

### v3

@@ 5.1 Results on Omni-Language Models
Gemini-1.5-Pro 42.91% 27.93% 26.09%
Reka-core-20240501 30.39% 23.12% 30.65%
Table 3: OLM Baselines Overall Results Grouped by Audio Type and Task Category. The accuracy numbers calculated by different audio types are at the upper table and the accuracy accross tasks categories are placed at the bottom table.
@@
Gemini-1.5-Pro 42.67% 42.26% 46.23%
Reka-core-20240501 31.52% 26.04% 33.02%
Accuracy ↑ Causal Inference (Temporal-)Spatial Entity Abstract Concept
Sub-class of QA Action & Activity Story Description Plot Inference Object Identification & Description Contextual & Environmental Identity & Relationship Text & Symbols Count & Quantity
MIO-Instruct (7B) 31.87% 21.30% 29.54% 60.66% 31.21% 31.25% 16.00% 6.67%
@@
Gemini-1.5-Pro 41.83% 30.87% 32.91% 62.56% 60.28% 31.25% 28.00% 13.33%
Reka-core-20240501 25.50% 24.78% 20.68% 49.76% 39.01% 28.12% 28.00% 6.67%
Breakdown Results.
We present the breakdown of the performance of open-source omni-language model baselines across different audio types and task categories in the OmniBench evaluation. The results reveal inconsistent performance patterns across audio types, with some models showing higher accuracy on sound events or music compared to speech. Across task categories, models tend to perform better on object identification and description tasks, while struggling with more reasoning tasks such as plot inference and story description, as illustrated by Table 3. This might be because visual entity recognition is an essential component for image captioning and other type of pre-training dataset. We observe Gemini provides significantly better results on the context/environment entities other than object entities. Furthermore, most of the models perform really bad on quantity & counting tasks. But scaling up of UnifiedIO model contributes a lot to this type of task. And scaling up from 1.1B to 3.2B benefits all senarios.
These findings highlight current limitations of OLMs in integrating information across modalities.
5.2 Textual Approximation on Images and Audios
Table 4: Results on Textual Audio Approximation Experiments. All the audios are represented in text transcript. The results are divided into groups of vision-language models and omni-models. We use the text transcript to approximate the audios in this setting. Boldface shows the best model performance, and underline shows the best open-source model.
@@ 5.2 Textual Approximation on Images and Audios
GPT4-o (0513) 57.62% 45.71% 42.21%
GPT4-o (0806) 51.14% 47.55% 31.44%
GPT4-o-mini 49.04% 39.23% 34.06%
Gemini-1.5-Pro 44.40% 22.50% 26.09%
Reka-core-20240501 46.58% 34.59% 30.65%
Claude-3.5-Sonnet 59.37% 33.54% 43.08%
GPT-4V-Preview 38.18% 41.24% 25.57%
GPT-4V-0409 33.36% 45.80% 32.75%
@@
As the absence of strong OLM baselines on OmniBench, we further introduce the text alternatives of images ($I^{\prime}$) and audios ($A^{\prime}$) to embrace more dual-modal MLLMs to analyze the current research progress in the field. The results using audio transcripts and image captions are put in Table 4, Table 5 correspondingly (results of $(I^{\prime},A^{\prime})$ setting at Table 6, Appendix A).
@@
LTU (7B) 23.29% 23.91% 23.12% 25.42% 20.00% 16.04%
Mu-LLaMA (7B) 1.58% 1.84% 1.84% 1.56% 1.13% 2.83%
MusiLingo-long-v1 13.66% 11.03% 9.02% 11.93% 13.96% 25.47%
Audio-SALMONN (13B) 34.76% 32.66% 33.36% 34.50% 29.43% 50.00%
Qwen-Audio-Chat (7B) 17.51% 16.64% 18.39% 14.66% 22.64% 25.47%
Qwen2-Audio-7B-Instruct 40.72% 35.20% 35.29% 40.60% 41.89% 38.68%
Audio-Flamingo (1.3B) 24.78% 23.82% 24.78% 26.98% 21.51% 16.98%
Gemini-1.5-Pro 38.62% 28.02% 21.02% 39.82% 33.96% 41.51%
Reka-core-20240501 29.42% 23.12% 26.27% 28.53% 29.43% 35.85%
@@
We select the UnifiedIO-2 series to conduct the textual approximation experiments due to their relatively robust performances in the vanilla evaluation setting suggested in Table 2. Compared with the vanilla setting, all three UnifiedIO-2 models show performance gains, averagely at $6.42\%$, in the audio replacement setting and average performance drops in the replaced-image ($1.87\%$) and both-repaced settings ($0.12\%$). This indicates the shortcoming of existing OLMs on modeling the audio on the one hand, and the potential noise in the generated image captions compared to the human-written audio transcripts on the other hand.
@@
In the setting of using text as the alternatives of audios and images, the VLMs show generally better results than ALMs (Table 4 vs Table 5) even compared with open-source model with similar model size. This could be caused by : 1) more available research resources have been put in VLMs to develop datasets and cross-modality alignment architectures, leading to higher instruction following rate and accuracy compared to ALMs; 2) the audio data are naturally harder (and hence more expensive) to annotate; and 3) audio typically has longer sequence tokens and requires more computational resource compared to text and image, making it harder to scale up. If $I^{\prime}$ and $A^{\prime}$ have the information loss ratio when converted from $I$ and $A$, it seems to be easier for the researchers to train the future omni-language models from exisiting VLMs rather than ALMs. Besides, we can observe Claude-3.5 and GPT-4o are generally the best two VLMs, significantly better compared to open-source VLMs. And Qwen2-audio and Gemini are the best two ALMs in speech and audio, and Audio-SALMONN is the best on music. Moreover, we can see significant difference on different type of audio, i.e., LTU and audio-flamingo are worse for music compare to speech and audio, while Qwen-audio which include music on pre-training provides better results on music compared to speech. And MusiLingo only use music for pre-training perform worse in speech and sudio.
Figure 5: The Performance Changes Brought By Textual Alternatives. The numbers in the cell suggest the accuracy change. (a) includes the UnifiedIO2 OLMs and the proprietary models supporting tri-modal inputs. (b) and (c) consists of VLMs and ALMs grouped with Gemini-1.5-Pro for comparison.
Pure Textual Evaluation.
The performance gaps brought by the replaced textual image and audio descriptions are in revealed in Figure 5. Notably, the majority of models demonstrate improved accuracy when processing textual representations of multimodal data compared to their performance on either image captions or audio transcripts alone. This suggests that these models show stronger reasoning capability when equipped with information from multiple textual sources rather than handling raw multimodal inputs. For instance, Qwen2-Audio-7B-Instruct shows a significant jump in accuracy from 39.05% (audio transcript only) and 39.67% (image caption only) to 47.02% when given both textual representations. Similarly, proprietary models like GPT4-o and Claude-3.5-Sonnet exhibit substantial gains, with GPT4-o (0513) achieving an impressive 60.60% accuracy in the pure textual setting.
6 Conclusion and Future Study
Ethics Statement
Our research on OmniBench and the development of multimodal language models raises several important ethical considerations:
  • Data Collection and Privacy: All image and audio data used in OmniBench was collected from public sources or created specifically for this research. We took care to remove any personally identifiable information. For human-recorded audio, participants provided informed consent and were compensated fairly for their time.
  • Potential Biases: We acknowledge that the dataset may contain inherent biases in terms of language, cultural representation, and types of scenarios depicted. We have made efforts to include diverse content, but further work is needed to fully characterize and mitigate these biases. Users of OmniBench should be aware of these limitations.
  • Responsible Disclosure: We will release OmniBench publicly to foster open research, but with appropriate use guidelines. The OmniInstruct dataset will be made available to researchers who agree to terms of responsible use.
We are committed to ongoing evaluation of the ethical implications of this work as the field of multimodal AI continues to advance rapidly.
Reproducibility Statement
We have made significant efforts to ensure the reproducibility of our work on OmniBench and the associated experiments:
  • Dataset: The complete OmniBench dataset, including all images, audio files, and question-answer pairs, will be made publicly available upon publication. Detailed information about the data collection process, annotation guidelines, and quality control measures are provided in section 3 and Appendix B.
  • Code: We have developed and will release a comprehensive codebase that includes: Scripts for data preprocessing and formatting; Implementation of all evaluation metrics; Code for running experiments.
  • Model Evaluation: For all baseline models evaluated, we provide detailed specifications. For proprietary models, we specify the exact API versions used and the dates of access.
  • Reproducibility Challenges: We acknowledge that exact reproduction of results for some proprietary models may be challenging due to potential API changes.
By providing these resources and detailed documentation, we aim to facilitate the reproduction of our results and encourage further research in this area. We welcome feedback from the community on any aspects that require additional clarification to ensure full reproducibility.
Appendix A More Experiment Results
@@
LTU (7B) 22.68% 24.17% 23.12%
Mu-LLaMA (7B) 2.28% 6.57% 1.84%
MusiLIngo-long-v1 (7B) 11.03% 10.51% 9.02%
Audio-SALMONN-13B 36.95% 34.41% 33.36%
Qwen-Audio-Chat 22.33% 24.69% 18.39%
Qwen2-Audio-7B-Instruct 46.15% 38.09% 35.29%
Audio-Flamingo (1.3B) 24.26% 23.73% 24.78%
@@
Cambrian-13B 43.08% 31.96% 29.16%
GPT4-o (0513) 60.51% 45.71% 37.92%
GPT4-o (0806) 53.77% 47.55% 29.51%
GPT4-o-mini 51.05% 49.04% 32.84%
Gemini-1.5-Pro 42.03% 22.50% 21.02%
Claude-3.5-Sonnet 56.83% 33.54% 39.05%
@@
Appendix B Dataset Development
B.1 Statistics for OmniInstruct Dataset
Table 7: The Statistics of Data Filtering in OmniInstruct. The table shows the number changes of question-answer pairs before and after filtering from each of the data sources.
Source Original Train Original Valid Remained Train Remained Valid
AVQA 40,182 16,798 4,491 (11.2%) 1,911 (11.4%)
Music-AVQA2.0 42,470 0 11 (0.03%) 0
MSRVTT-QA 140,554 11,143 80,078 (57.0%) 6,479 (58.1%)
Total 233,206 27,941 84,580 8,390
As demonstrate in Table 7, most of the samples in the dataset are in low quality and therefore abandoned, and only 93k of samples remain for Omni-modality SFT training. This is reasonable because most of the questions are generated from templates, and the image may not sampled from the most relevant part of the questions and, therefore hot high in quality.
B.2 Prompt for Quality Control on OmniInstruct Dataset
Figure 6: The Prompt for OmniInstruct Dataset Filtering.
B.3 Diversity of Music Audios of OmniBenchmark
The music subset of our benchmark reflects a rich diversity of musical traditions, spanning a wide range of genres, styles, and cultural contexts. It encompasses Western classical symphonies, jazz chamber music, and avant-garde compositions alongside popular music from China, England, and France. Traditional forms like Kunqu opera and modern experimental pieces are represented, as well as instrumental music from regions such as India, the Arab world, Africa, and Japan. The benchmark also includes famous film soundtracks with various thematic elements and Asian folk oral traditions, such as chanting, drumming, and Humai. This eclectic collection, enriched by unique instances like famous concert spoofs and iconic YouTube parodies, ensures that each question offers a distinct challenge, showcasing the nuanced intricacies and breadth of global music heritage.